# 1] organization=School of Computer Sciences , addressline=Universiti Sains Malaysia, city=11800 USM Penang, country=Malaysia

### 23 Dec 2023 v5

@@ Abstract
Despite the remarkable results of deep learning in breast cancer histopathology image classification, challenges such as data imbalance and interpretability still exist and require cross-domain knowledge and collaboration among medical experts. This study proposes a breast cancer classification method using a dual-activated lightweight attention ResNet50 model, effectively addressing data imbalance and interpretability challenges. The model fuses a pre-trained deep ResNet50 and a lightweight attention mechanism to accomplish classification by embedding an attention module in layer 4 of ResNet50 and adding two fully connected layers. The fully connected network design employs LeakyReLU and ReLU activation functions.
The model outperforms SEResNet50, DensNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50 models regarding precision, accuracy, recall, F1 score, and GMean, especially in the application performance on the BreakHis dataset. In particular, the model demonstrates significant robustness and broad applicability when dealing with the unbalanced breast cancer dataset. The model has been evaluated on histopathology images at magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. The study comprehensively assessed the model’s performance. In the later stages of training, the validated losses and accuracies change minimally, showing that the model avoids overfitting and exhibits good generalization ability. This model exhibited the fastest convergence in all laboratory experiments, even though its parameters are not the smallest. This highlights the model’s efficacy as a lightweight attention framework, showcasing its efficiency in achieving rapid convergence without compromising performance.
keywords:
@@
1 Introduction
@@
Breast cancer, particularly among women, has emerged as one of the leading causes of fatal diseases, with its incidence on a continuous rise, eliciting worldwide health concerns [1, 2]. Early and accurate diagnosis and treatment are crucial in reducing the mortality rate of breast cancer. However, medical practices grapple with several obstacles, including inadequate healthcare resources and human error in interpretation, which hampers accurate and prompt detection and diagnosis of breast cancer. In this context, intelligent assisted diagnosis technology has become an innovative solution in breast cancer. Recent advancements in deep learning have had a transformative impact on breast cancer diagnosis. Deep learning theory has not only surpassed the performance limitations of traditional classification and recognition algorithms but has also integrated the entire feature extraction, feature selection, and classification process into a single step. Deep learning has simplified the training process by automatically extracting and classifying features through neural networks, eliminating redundancy and improving efficiency. These developments have revolutionized breast cancer diagnosis by improving accuracy and streamlining workflows [3]. This profound learning advancement has revolutionized breast cancer diagnosis [4, 5, 6, 7]. One key innovation in this area is developing and applying Convolutional Neural Networks (CNNs), a specific class of deep learning models designed to process data with a grid-like structure, such as medical images. The concept of CNN was first introduced back in 1998 [8]. The concept describes a method for efficiently classifying image data. CNNs gradually extract abstract features of an image through multi-level convolutional and pooling operations, allowing the network to extract necessary features from different categories of images so that the different categories can be classified accurately based on those features. CNNs have three main layers: convolutional, pooling, and fully connected. These layers work together to make CNNs a powerful tool for classifying medical images. The convolutional layer captures the image’s local features, the pooling layer reduces data dimensions while preserving significant information, and the fully connected layer maps these extracted features to various output classes, As shown in Fig. 1. However, despite the significant achievements of deep learning in medical image classification, conventional CNN architectures struggle with unbalanced medical image data. To tackle this issue, In this paper, a lightweight attention mechanism based on ResNet50 [9] is introduced to address the challenge of unbalanced medical image data in conventional CNN architectures. This novel approach effectively handles unstable category distributions and accurately detects malignant breast cancer in medical images. The critical contributions of the work can be summarized as follows:
@@
(b) Integrating the attention mechanism module and ResNet50 to form a feature-rich model, the model achieves more efficient feature extraction based on fusing the two techniques. Specifically, in the fully connected network part of the attention mechanism module, we employed LeakyReLU [10] and ReLU [11] activation functions. Leaky-ReLU was chosen to mitigate the gradient vanishing problem, while ReLU was used to enhance the image feature representation, thus better capturing multi-scale spatial patterns.
@@
Fig. 1: Overview of Convolutional Neural Network
2 Related works
@@ 2.1 CNN
The Inception model is a neural network architecture introduced by Google, which aims to solve the computation problem and several parameters in neural networks and is suitable for image classification and related tasks. The hybrid model based on CNN-Inception-V4 was used by Nazir et al. [18]. It has excellent accuracy and sensitivity, but the model’s specificity is a little inferior, and it does not consider the processing computational efficiency and resource consumption issues.
@@
2.2 Transfer learning
Transfer learning applies pre-trained models from one image domain to another, effectively bridging the gap between different image domains. In the context of breast cancer image classification, this pre-trained model approach involves fine-tuning CNN models, which were initially trained on a comprehensive large-scale image data set, and then fine-tuning it to adapt to the data set specifically used for breast cancer detection. Through this step, the model is optimized to better meet the specific needs of breast cancer classification. By using the inherent ability of the pre-trained model in generalization and feature extraction, this strategy significantly improves the efficiency of training and the accuracy of classification performance [23, 24, 25, 26, 27, 28, 29].
@@ 2.2 Transfer learning
In conclusion, considering the exceptional robustness of ResNet50 in transfer learning, this study has chosen to further optimize and enhance it. However, we face specific challenges when dealing with imbalanced breast cancer datasets. Firstly, the imbalance in the dataset tends to bias the model towards categories with a larger number of instances, potentially leading to a significant reduction in the detection rate of key but less frequent categories, especially in early-stage breast cancer cases, thereby increasing the risk of missed diagnoses. Secondly, given the limited sample size of minority categories, there is a risk of overfitting in these samples, affecting the model’s generalization ability for novel and previously unseen minority class samples. Therefore, the goal of this study is to overcome these challenges by meticulously optimizing the application of ResNet50 in transfer learning, thereby enhancing the model’s overall performance in dealing with imbalanced breast cancer data.
2.3 Attention mechanism
@@
In the field of medical imaging, the attention mechanism [32] in deep learning has proven to be a groundbreaking technique. It enables models to assign varying levels of importance to different regions within medical images, such as MRI scans, X-rays, CT scans, and histopathology images [33]. This targeted focus allows the models to prioritize and analyze the most pertinent areas for diagnosing diseases, identifying anomalies, or tracking the progression of medical conditions. By concentrating on crucial features within these images, the attention mechanism greatly enhances the accuracy and efficiency of automated medical image analysis, aiding in early detection and improved patient outcomes. In the context of breast cancer image classification, introducing the attention mechanism approach can prompt the model to learn critical and discriminative features in breast cancer images that are crucial for accurately distinguishing between malignant and benign tissues. This technique can improve the model’s performance while reducing the reliance on global image feature extraction and improving the accuracy of identifying critical breast cancer regions [34].
@@
Researchers have proposed various breast cancer segmentation and classification systems incorporating attention mechanisms. For instance, Vardhan et al. [35] developed a breast cancer segmentation system that uses convolutional neural networks with interpretable artificial intelligence and the attention mechanism to improve breast cancer detection and classification accuracy and speed. Xu et al. [36] designed a system for breast cancer segmentation that uses a modified DenseNet architecture with an attention mechanism to optimize the convolutional neural network on the BreakHis dataset for high-precision classification of breast cancer. The method was tested on the BreakHis dataset, and the method’s performance was evaluated using accuracy metrics for both binary and multi-classification. Wang et al. [37] proposed a breast cancer classification model that combines deep migration learning and visual attention mechanisms to improve the accuracy and efficiency of real-world clinical diagnosis. This deep-migration-based attention mechanisms approach uses migration learning to overcome the overfitting problem and combines the attention module to focus on characterizing breast lesions. Wang et al.’s work [37] achieves high-precision results that outperform other state-of-the-art breast image classification algorithms. This section summarizes recent scholarly research and exploration in breast cancer classification, focusing on the use of the attention mechanism. The insights gleaned indicate that applying the attention mechanism in breast cancer image classification methods can enhance the accuracy of crucial region identification and boost the interpretability of the model simultaneously. Still, the difficulty of training and parameter tuning is also increased accordingly. In this study, a lightweight design of the attention module is proposed, which incorporates adaptive average pooling, a two-layer fully connected network, and the dual activation function of LeakyReLU and ReLU to efficiently extract and enhance the critical information in the image through item-by-item multiplication (element-wise multiplication of individual pixels and features within the images)of features images, aiming to improve performance and reduce computational consumption.
3 Methods
3.1 Transfer ResNet50
@@
The ResNet50 [9] model has a relatively complex structure, including a 50-layer deep network structure. The specific structure is shown in Table 1. The primary module is made up of numerous residual blocks. Each block is equipped with multiple convolutional layers and distinctive identity maps. Uniquely, the ResNet50 model introduces the concept of residual learning. This innovation solves the problem of vanishing gradients in deep networks, thereby opening up the possibility of building deeper neural networks than ever before. This design breakthrough allowed for achieving exceptional performance. However, considering the efficient utilization of data and computing resources, the transfer learning strategy is adopted in this paper based on the pre-trained ResNet50 model for model building. Through transfer learning, full benefit can be obtained from the high-level feature representations learned by ResNet50 on large-scale datasets in medical image classification tasks. This allows for swiftly attaining superior classification outcomes, even with constrained data and computing capacities.
Table 1: The ResNet50 Model
@@
Note: The term "1000-d FC" in the pooling layer refers to a fully-connected layer consisting of 1000 neurons.
3.2 Dual-activated lightweight attention ResNet50
The paper proposes a dual-activation lightweight attention ResNet50 model method. Initially, a pre-trained ResNet50 is utilized as the foundational architecture. Building upon this, an innovative dual-activation lightweight attention module is introduced, along with a newly designed fully connected layer tailored for specific tasks. This attention module is integrated into the fourth residual block of ResNet50, characterized by the amalgamation of adaptive average pooling and a two-layer fully connected network, employing LeakyReLU[10] and ReLU [11] as its dual activation functions. Additionally, the newly added fully connected layer is positioned at the end of the modified ResNet50 model, primarily aimed at addressing the complexities of binary classification problems.
In the dual-activation lightweight attention module proposed in this paper, adopt the LeakyReLU activation function [10] for the initial fully connected layer, leveraging its competence in allowing for a slight gradient flow under zero. This is done to mitigate the issue of vanishing gradients, thereby enhancing the training efficacy of deep neural networks. The ReLU activation function [11] is chosen for the second fully connected network to further improve the model performance and image feature representation.
Combining these two activation functions greatly enhances the model’s expressive power and capacity to learn intricate image features, making the model more suitable for image classification tasks. In the experiments, this design improves accuracy and performance while maintaining the convergence speed of the model. The flowchart and visualization of the "Dual Activation Lightweight Attention ResNet50 Module" can be comprehended through the displayed in Fig. 2 and Fig. 3, respectively, and the overall technical framework is shown in Fig. 4.
Fig. 2: Flowchart for dual-activated lightweight attention ResNet50 model
Fig. 3: Visualization of dual-activated lightweight attention ResNet50 model
Fig. 4: The framework of dual-activated lightweight attention ResNet50 model
4 Experiment
4.1 BreakHis Dataset

### 14 Feb 2024 v6

@@ Abstract
Automatic breast cancer classification technology in histopathology images plays a crucial role in accurate diagnosis and treatment planning. Recently, histopathology image classification approaches based on ResNet architecture have been widely proposed as they vastly improve the classification accuracy using skip connection to prevent vanishing gradient issues by combining low-level and high-level feature information. However, traditional ResNet architecture suffers from challenges such as data imbalance and interpretability, which still require cross-domain knowledge and collaboration among medical experts. This study proposes a new breast cancer classification method using a Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) approach to address data imbalance and interpretability challenges effectively. The proposed approach combines a pre-trained deep ResNet50 method and a lightweight attention mechanism to accomplish the classification by embedding an attention module in layer 4 of ResNet50 and adding two fully connected layers based on LeakyReLU and ReLU activation functions to enhance the ability of feature learning. The proposed DALAResNet50 approach has been evaluated on breast cancer histopathology images derived from BreakHis Datase at magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. The DALAResNet50 approach was comprehensively assessed with popular deep learning approaches such as SEResNet50, DensNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50 models. The reported results of DALAResNet50 outperform the compared approaches regarding precision, accuracy, recall, F1 score, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and unbalanced breast cancer datasets.
Index Terms:
Lightweight attention mechanism, Convolutional Neural Networks, breast cancer classification.
@@
=-21pt
Figure 1: Overview of Convolutional Neural Network
I Introduction
@@
Breast cancer, particularly among women, has emerged as one of the leading causes of fatal diseases, with its incidence on a continuous rise, eliciting worldwide health concerns [1, 2]. Early and accurate diagnosis and treatment are crucial in reducing the mortality rate of breast cancer. However, medical practices grapple with several obstacles, including inadequate healthcare resources and human error in interpretation, which hampers accurate and prompt detection and diagnosis of breast cancer. In this context, intelligent assisted diagnosis technology has become an innovative solution in breast cancer. Recent advancements in deep learning have had a transformative impact on breast cancer diagnosis. Deep learning theory has not only surpassed the performance limitations of traditional classification and recognition algorithms but has also integrated the entire feature extraction, feature selection, and classification process into a single step. Deep learning has simplified the training process by automatically extracting and classifying features through neural networks, eliminating redundancy and improving efficiency. These developments have revolutionized breast cancer diagnosis by improving accuracy and streamlining workflows [3]. This profound learning advancement has revolutionized breast cancer diagnosis [4, 5, 6, 7]. One key innovation in this area is developing and applying Convolutional Neural Networks (CNNs), a specific class of deep learning models designed to process data, such as medical images, with a grid-like structure. The concept of CNN was first introduced back in 1998 [8]. The concept describes a method for efficiently classifying image data. CNNs gradually extract abstract features of an image through multi-level convolutional and pooling operations, allowing the network to remove necessary features from different categories of pictures so that the other categories can be classified accurately based on those features. CNNs have three main layers: convolutional, pooling, and fully connected. These layers work together to make CNNs a powerful tool for classifying medical images. The convolutional layer captures the image’s local features, the pooling layer reduces data dimensions while preserving significant information, and the fully connected layer maps these extracted features to various output classes, As shown in Fig. 1. However, despite the considerable achievements of deep learning in medical image classification, conventional CNN architectures struggle with unbalanced medical image data. To tackle this issue, In this paper, a lightweight attention mechanism based on ResNet50 [9] is introduced to address the challenge of unbalanced medical image data in conventional CNN architectures. This novel approach effectively handles unstable category distributions and accurately detects malignant breast cancer in medical images. The critical contributions of the work can be summarized as follows:
@@
(b) Integrating the attention mechanism module and ResNet50 to form a feature-rich model, the model achieves more efficient feature extraction based on fusing the two techniques. Specifically, in the fully connected network component of the attention mechanism module, LeakyReLU [10] and ReLU [11] activation functions were utilized. LeakyReLU was selected for its ability to mitigate the gradient vanishing problem. In contrast, ReLU was employed to enhance the representation of image features. This choice was instrumental in better capturing multi-scale spatial patterns.
@@
II Related works
@@ II-A CNN
The Inception model is a neural network architecture introduced by Google that aims to solve the computation problem and several parameters in neural networks and is suitable for image classification and related tasks. The hybrid model based on CNN-Inception-V4 was used by Nazir et al. [18]. It has excellent accuracy and sensitivity, but the model’s specificity is a little inferior, and it does not consider the processing computational efficiency and resource consumption issues.
@@
II-B Transfer learning
Transfer learning applies pre-trained models from one image domain to another, bridging the gap between different image domains. In the context of breast cancer image classification, this pre-trained model approach involves fine-tuning CNN models, which were initially trained on a comprehensive large-scale image data set, and then fine-tuning it to adapt to the data set used explicitly for breast cancer detection. This step optimizes the model better to meet the specific needs of breast cancer classification. By using the inherent ability of the pre-trained model in generalization and feature extraction, this strategy significantly improves the efficiency of training and the accuracy of classification performance [23, 24, 25, 26, 27, 28, 29].
@@ II-B Transfer learning
In conclusion, this study has focused on further optimization and enhancement due to the exceptional robustness of ResNet50 in transfer learning. Nevertheless, the investigation encounters specific challenges when dealing with imbalanced breast cancer datasets. Firstly, the imbalance in the dataset tends to bias the model towards categories with a larger number of instances, potentially leading to a significant reduction in the detection rate of key but less frequent categories, especially in early-stage breast cancer cases, thereby increasing the risk of missed diagnoses. Secondly, given the limited sample size of minority categories, there is a risk of overfitting in these samples, affecting the model’s generalization ability for novel and previously unseen minority class samples. Therefore, this study aims to overcome these challenges by meticulously optimizing the application of ResNet50 in transfer learning, thereby enhancing the model’s overall performance in dealing with imbalanced breast cancer data.
II-C Attention mechanism
@@
In medical imaging, the attention mechanism [32] in deep learning has proven to be a groundbreaking technique. It enables models to assign varying levels of importance to different regions within medical images, such as MRI scans, X-rays, CT scans, and histopathology images [33]. This targeted focus allows the models to prioritize and analyze the most pertinent areas for diagnosing diseases, identifying anomalies, or tracking the progression of medical conditions. By concentrating on crucial features within these images, the attention mechanism dramatically enhances the accuracy and efficiency of automated medical image analysis, aiding in early detection and improved patient outcomes. In the context of breast cancer image classification, introducing the attention mechanism approach can prompt the model to learn critical and discriminative features in breast cancer images crucial for accurately distinguishing between malignant and benign tissues. This technique can improve the model’s performance while reducing the reliance on global image feature extraction and improving the accuracy of identifying critical breast cancer regions [34].
@@
Researchers have proposed various breast cancer segmentation and classification systems incorporating attention mechanisms. For instance, Vardhan et al. [35] developed a breast cancer segmentation system that uses convolutional neural networks with interpretable artificial intelligence and the attention mechanism to improve breast cancer detection and classification accuracy and speed. Xu et al. [36] designed a system for breast cancer segmentation that uses a modified DenseNet architecture with an attention mechanism to optimize the convolutional neural network on the BreakHis dataset for high-precision classification of breast cancer. The method was tested on the BreakHis dataset, and the method’s performance was evaluated using accuracy metrics for both binary and multi-classification. Wang et al. [37] proposed a breast cancer classification model that combines deep migration learning and visual attention mechanisms to improve the accuracy and efficiency of real-world clinical diagnosis. This deep-migration-based attention mechanisms approach uses migration learning to overcome the overfitting problem and combines the attention module to focus on characterizing breast lesions. Wang et al.’s work [37] achieves high-precision results that outperform other state-of-the-art breast image classification algorithms. This section summarizes recent scholarly research and exploration in breast cancer classification, focusing on the use of the attention mechanism. The insights gleaned indicate that applying the attention mechanism in breast cancer images and classification methods can simultaneously enhance the accuracy of crucial region identification and boost the interpretability of the model. However, this also correspondingly increases the following difficulties:
II-C1 Limitations in global feature extraction
Although the attention mechanism is designed to enhance model performance by focusing on critical parts of the image, thereby playing a positive role in strengthening the model’s ability to process global information, it may still face challenges in some implementations or specific application scenarios, unable to fully capture the international contextual details of the entire image.
II-C2 Insufficiency in feature priority precision
Lacking a sophisticated method to discern the significance of diverse features, attention mechanisms might uniformly weigh all channel information. This approach risks overlooking vital features essential for precise classification, undermining the model’s ability to highlight and utilize discriminative patterns within the data effectively.
II-C3 Inefficiency of Activation Functions
The reliance on ReLU within attention mechanisms can lead to the ”dying ReLU” issue, wherein neurons fail to activate over specific data ranges, stalling the learning progression. Such constraints significantly diminish the model’s adaptability during the training phase, adversely affecting the dynamics and overall efficacy of the attention-enhanced learning process.
II-C4 Lack of feature adaptability
Attention mechanisms may exhibit rigidity in adjusting and enhancing features non-linearly according to the unique characteristics of each input. This inflexibility can constrain the model’s ability to adapt and optimally leverage the diverse information within the data, thereby reducing the overall efficiency in feature utilization.
This study designed an innovative lightweight attention module to overcome the identified difficulties. This design incorporates adaptive average pooling, a two-layer fully connected network for channel recalibration, dual activation functions of LeakyReLU [10] and ReLU [11], and element-wise multiplication. Adaptive average pooling tackles the limitations in global feature extraction, whereas the two-layer fully connected network for channel recalibration optimizes feature priority precision. The combination of LeakyReLU and ReLU effectively addresses the inefficiencies of a single activation function, enhancing activation function efficiency, and element-wise multiplication improves feature adaptability. These pioneering approaches synergistically aim to significantly elevate model performance while mitigating computational resource requirements, providing a more efficient and adaptable solution for integrating attention mechanisms into breast cancer image classification models.
III Methods
III-A Transfer ResNet50
@@
The ResNet50 [9] model has a relatively complex structure, including a 50-layer deep network structure. The specific structure is shown in Table I. The primary module is made up of numerous residual blocks. Each block is equipped with multiple convolutional layers and distinctive identity maps. Uniquely, the ResNet50 model introduces the concept of residual learning. This innovation solves the problem of vanishing gradients in deep networks, thereby opening up the possibility of building deeper neural networks than ever before. This design breakthrough allowed for achieving exceptional performance. However, considering the efficient utilization of data and computing resources, the transfer learning strategy is adopted in this paper based on the pre-trained ResNet50 model for model building. Through transfer learning, full benefit can be obtained from the high-level feature representations learned by ResNet50 on large-scale datasets in medical image classification tasks. This allows for swiftly attaining superior classification outcomes, even with constrained data and computing capacities.
TABLE I: The ResNet50 Model
@@
Note: The term ”1000-d FC” in the pooling layer refers to a fully-connected layer consisting of 1000 neurons.
TABLE II: Comparison between DALAResNet50 model and standard ResNet model
Feature
DALAResNet50 model
Standard ResNet model
Adaptive Average Pooling
Adaptive average pooling addresses ”limitations in global feature extraction” by enhancing the model’s capacity to process global contextual information, significantly improving performance in tasks requiring comprehensive image understanding.
Not available
Channel Recalibration
A two-layer fully connected network for channel recalibration optimizes ”feature priority precision,” enhancing the model’s ability to identify and focus on the most relevant features for accurate classification.
Not available
Activation Functions
The dual activation functions of LeakyReLU and ReLU address the ”inefficiency of activation functions.” By incorporating LeakyReLU, the network avoids the dying ReLU problem, as it allows a small, non-zero gradient when the unit is not active, ensuring continuous learning dynamics and preventing neurons from becoming inactive across all network layers. This combination maintains the model’s learning capability, potentially enhancing the training process.
ReLU only
Element-wise multiplication
Element-wise multiplication, as denoted by ’Scale’ in Fig. 3, enhances ”Feature Adaptability” by allowing dynamic, non-linear adjustment of features according to a specific input, improving the network’s flexibility and utilization of information.
Not available
Fig. 2: Flowchart for standard ResNet model
Fig. 3: Flowchart for DALAResNet50 model
Fig. 4: Visualization of DALAResNet50 model
Fig. 5: The framework of DALAResNet50 model
III-B Dual-activated lightweight attention ResNet50
A new approach for automatic breast cancer classification is proposed in this research using a new version of the ResNet50 technique called Dual-activated lightweight attention ResNet50 (DALAResNet50). The DALAResNet50 approach incorporates ResNet50, Global pooling, channel recalibration, Dual Activation functions, and the newly designed fully connected layer into a unified approach to overcome the drawbacks of the traditional ResNet50 model. The pre-trained ResNet50 is first utilized as a baseline architecture. Subsequently, the attention module is integrated into the fourth residual block of ResNet50 to pay attention to the small-scale nuclei cells of breast cancer. The attention mechanism improves the ability of feature learning by extracting the silent semantic information to overcome the class imbalance issue. The proposed attention module is characterized by the amalgamation of adaptive global pooling and a two-layer fully connected network, employing LeakyReLU [10] and ReLU [11] as dual activation functions. The traditional ResNet50 uses conventional pooling layers to reduce the spatial dimensions of the feature maps while retaining important information. However, the conventional pooling layers operate over local regions of the input feature maps, reducing their spatial dimensions but discarding spatial information, making the ResNet50 sensitive to spatial translations of the input feature map and increasing the number of parameters in the network. Therefore, this research proposed average-based global pooling that calculates the average value of each feature map across all spatial locations, effectively summarizing each feature’s presence throughout the entire feature map. The main advantage of global pooling is that it provides a fixed-size representation of the whole feature map, making the ResNet50 less sensitive to spatial translations of the input and reducing the total number of parameters in the network. Additionally, the DALAResNet50 adopts the LeakyReLU activation function [10] for the initial fully connected layer, leveraging its competence in allowing for a slight gradient flow under zero. This is done to mitigate the issue of vanishing gradients, thereby enhancing the training efficacy of deep neural networks. The ReLU activation function [11] is further chosen for the second fully connected network to improve the model performance and image feature representation. Combining these two activation functions enhances the model’s expressive power and capacity to learn intricate image features, making the model more suitable for breast cancer image classification tasks. Finally, the newly added fully connected layer is positioned at the end of the DALAResNet50 approach to address the complexity of binary classification problems.
Fig. 2 and Fig. 3 illustrate the flowcharts of the standard ResNet approach and the DALAResNet50 approach, respectively. These flowcharts highlight the newly added features in the DALAResNet50. A comprehensive comparative analysis of these differences is detailed in Table 2. The visualization of the DALAResNet50 approach and the entire technical framework are depicted in Fig. 4 and Fig. 5.
IV Experiment
IV-A BreakHis Dataset